OpenStudy (anonymous):

Which equation is shown on the graph? A. y = x + 2 B. y = 2x C. y = x + 1 D. y = x – 2

OpenStudy (anonymous):

@chiko_1278

OpenStudy (anonymous):

graph ?

OpenStudy (anonymous):

yeah....but what graph i dont see any graph in here

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

this is the graph

OpenStudy (anonymous):

i think is A

OpenStudy (anonymous):

or B

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

it think its a

OpenStudy (anonymous):

yea cuz 2/2=1 and it crosses the y axis at 2

OpenStudy (anonymous):

thanks @tylermcmullen23 and @Chiko_1278

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

need anymore help?

OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 3x – 4 A. (11, 5) and (2, 2) B. (5, 11) and (3, 2) C. (5, 11) and (2, 3) D. (5, 11) and (2, 2)

OpenStudy (anonymous):

its showing as boxes for me?

OpenStudy (anonymous):

is it just y= 3x+4?

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

@Chiko_1278

OpenStudy (anonymous):

@openstudyrocks5*

OpenStudy (anonymous):

what do you thnik it is

OpenStudy (openstudyrocks5*):

Just so you know their should only be 1 question in a question :P......

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

:L

OpenStudy (openstudyrocks5*):

Wow I honestly don't know... Hold on 1 sec.

OpenStudy (anonymous):

i think d

OpenStudy (anonymous):

i think d aswell

OpenStudy (anonymous):

i checked using a graph. it should be D

OpenStudy (openstudyrocks5*):

For the equation.. Does this help?

OpenStudy (anonymous):

the equation is -4 though not +4

OpenStudy (anonymous):

no but that was kind of you

OpenStudy (openstudyrocks5*):

It keeps saying +

OpenStudy (openstudyrocks5*):

Here. http://www.cymath.com/

OpenStudy (openstudyrocks5*):

@1DISAMAZING2345

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=3x−4 Math Keyboard ▼ 1 Add 4 to both sides y+4=3x 2 Divide both sides by 3 y+43=x 3 Switch sides x=y+43 Done